# Quandles as pre-Lie skew braces, set-theoretic Hopf algebras &\&& universal ${\cal R}$-matrices

### v1

@@ Abstract.
We present connections between left non-degenerate solutions of set-theoretic Yang-Baxter equation and left shelves using certain maps called Drinfel’d homomorphisms. We further generalise the notion of affine quandle, by using heap endomorphisms and metahomomorphisms, and identify the Yang-Baxter algebra for solutions of the braid equation associated to a given quandle. We introduce the notion of the pre-Lie skew brace and identify certain affine quandles that give rise to pre-Lie skew braces. Generalisations of the braiding of a group, associated to set-theoretic solutions of the braid equation is also presented. These generalized structures encode part of the underlying Hopf algebra. Indeed, we also introduce the quasi-triangular Hopf algebras and the universal ${\cal R}$-matrices for quandle algebras and for set-theoretic Yang-Baxter algebras. In fact, we obtain the universal ${\cal R}$-matrix for the set-theoretic Yang-Baxter algebras after identifying the associated admissible Drinfel’d twist. Generic set-theoretic solutions coming from heap endomorphisms are also identified.
@@ 1. Introduction
Following [20], given a set $X$, a map $r:X\times X\to X\times X$ is said to be a set-theoretic solution of the Yang-Baxter equation, if $r$ satisfies the braid identity
@@
The key outcomes of this investigation are presented in Sections 3-5. In these sections we generalize the notion of affine quandle, by using heap endomorphisms and metahomomorphisms, and identify the Yang-Baxter algebra for solutions of the braid equation associated to a given quandle. Motivated by the notion of pre-Lie algebras (also studied under the name chronological algebras) [1, 26, 41] (see also [4, 33] for a recent reviews) we introduce a novel algebraic structure called pre-Lie skew brace to describe the underlying algebraic structures associated to certain set-theoretic solutions of the braid equation. In fact, we identify families of affine quandles that produce pre-Lie skew braces. We then consider the linearized version of the braid equation and we introduce the notion of the quasi-triangular Hopf algerbas associated to set-theoretic solutions; the quandle type universal ${\cal R}-$matrix is also derived. By identifying a suitable admissible Drinfel’d twist [21, 22] we are able to extract the general set-theoretic universal ${\cal R}-$matrix.
More precisely, the outline of the paper is as follows.
@@
We start with Section 2 which is divided into two subsections. The first subsection contains preliminaries on shelves. In the second subsection, in 2.9, we introduce the notion of a Drinfel’d homomorphism between set-theoretic solutions of the Yang-Baxter equation, based on the well known idea of Drinfel’d twists. When one considers a Drinfel’d automorphism, one gets a set-theoretic analogue of admissible twists of quasitriangular Hopf-algebras. We show that every left non-degenerate solution is Drinfel’d isomorphic to a solution given by a shelf, see Lemma 2.12. In fact, by 2.15, every left non-degenerate solution can be acquired from shelf solution by finding special automorphisms of the shelf itself, which we name twists, see 2.14. In particular, bijective solutions correspond to racks. We show that two homomorphic shelves with twists have D-homomorphic solutions, and bijectivity of a homomorphism of shelves implies D-isomorphism of solutions. We conclude Section 2 with examples of solutions given by shelves and twists on them.
@@
Section 3 is divided into three subsections. In the first subsection, we present a definition of a Yang-Baxter algebra $(X,m)$ associated with a given solution $(X,r)$ as a binary operation $m:X\times X\to X$ such that $mr=m,$ i.e., by Yang-Baxter algebra in 3.1; specifically, we understand an algebra in the sense of a universal algebra that is invariant under the entanglement of a particular set-theoretic solution of the Yang-Baxter equation. We show that for any set endowed with a left non-degenerate solution, the Yang-Baxter algebra exists, see Lemma 3.4. We conclude the first subsection with examples of Yang-Baxter algebras for solutions given by skew braces and near braces. In the second subsection, we introduce three generalisations of affine quandles to the non-abelian case and study their Yang-Baxter algebras, see examples 3.12,3.13 and 3.14. In particular, endofunctions called metahomomorphisms and heap endomorphisms allow us to define such generalised quandle. Moreover, if the underlying group of those quandles is abelian and metahomomorphisms are group endomorphisms, all the examples collapse to the definition of an affine quandle. In addition, every heap endomorphism gives rise to quandle given by a metahomomorphism, see Lemma 3.16; nevertheless, the Yang-Baxter algebra operation obtained by a heap endomorphism is not necessarily given by a metahomomorphism, see 3.11 and Lemma 3.20. In the third and last subsection, we introduce a notion of the right pre-Lie skew brace, see 3.23. In particular, every pre-Lie ring is a right pre-Lie brace. Next, we show that Yang-Baxter algebras from examples 3.13 and 3.14 admit a pre-Lie skew brace structure, see 3.28. If an underlying group of a pre-Lie skew brace is abelian, one can use tools from [6] and [7] to acquire a pre-Lie ring. We conclude the subsection with 3.34 in which we show that Lie rings associated with a Yang-Baxter algebra given by affine quandle have a zero Lie bracket.
@@
In Section 4, we present a generalised version of the braiding of a group, with some examples that have already appeared in [19, 18], see 4.1. We generalise it beyond the group structure to be compatible with Yang-Baxter algebras. We conclude the section with examples of such braidings, see examples 4.6,4.7,4.8 and 4.9. This section serves as a precursor of the last section given that the generalized braided structures encapsulate part of the associated underlying Hopf algebras, which is introduced next.
@@
In Section 5, which is divided in two subsections, we shift our focus from set-theoretic solutions to linearized versions and Hopf algebras. Specifically, in the first subsection, introduce the Yang-Baxter algebras of quandle type and set-theoretic solutions as quasitriangular (quasi)-Hopf algebras [21, 22]. Relevant earlier works in [16, 17, 18] also in [23] and [2] in connection with pointed Hopf algebras, and the bialgebras associated to racks are studied in [15, 31, 3]. We first introduce the quandle Hopf algebra and we then systematically construct the associated universal ${\cal R}$-matrix. In the second subsection, we suitably extend the quandle algebra and present the set-theoretic Yang-Baxter algebra, which is also a Hopf algebra. By means of a a suitable admissible universal Drinfel’d twist we construct the universal-set theoretic ${\cal R}-$matrix. We note that the linearized version of the Yang-Baxter equation allows the investigation of the full Hopf algebras and not only the study of the subset consisting of group-like elements that naturally appear in the set-theoretic frame. Fundamental representations of the aforementioned algebras are also considered leading to the quandle and set-theoretic solutions of the Yang-Baxter equation. Generalized novel set-theoretic solutions coming from heap endomorphisms and being compatible with the universal set-theoretic Yang-Baxter algebra are also introduced.
@@ 2.1. Left shelves
Let $X$ be a non-empty set and  normal-$\triangleright$  a binary operation on $X$. Then, the pair $\left(X,\,\triangleright\right)$ is said to be a left shelf if  normal-$\triangleright$  is left self-distributive, namely, the identity
@@
If $\left(X,\,\triangleright\right)$ and $\left(Y,\,\blacktriangleright\right)$ are left shelves, a map $f:X\to Y$ is said to be a shelf homomorphism if $f\left(a\triangleright b\right)=f\left(a\right)\blacktriangleright f\left(b\right)$, for all $a,b\in X$, or, equivalently, $fL_{a}\left(b\right)=\bar{L}_{f\left(a\right)}f\left(b\right)$, for all $a,b\in X$, where $\bar{L}_{u}\left(v\right)=u\blacktriangleright v$, for all $u,v\in Y$. We denote the set of all endomorphisms of $\left(X,\,\triangleright\right)$  by $\mathrm{End}(X,\,\triangleright)$.
@@ 2.2. Left-non degenerate solutions and left shelves
The converse of the previous fact does not hold
@@
Given a set $X$, a map $r:X\times X\to X\times X$ satisfies (1.1) if and only if the following three conditions are satisfied for all $a,b,c\in X$
@@
$=\varphi^{-1}_{\varphi_{\tau_{b}\left(a\right)}\left(c\right)}L_{\varphi^{-1}_{\varphi_{a}\left(b\right)}\varphi_{a}\varphi_{b}\left(c\right)}\varphi^{-1}_{\varphi_{a}\left(b\right)}L_{\varphi_{a}\left(b\right)}\left(a\right)$ $\varphi_{a}\in End\left(X,\,\triangleright\right)$
@@
As a consequence of 2.16 and [11, Corollary 6] we can state the following result.
@@
If $\left(X,\,\triangleright\right)$ is a left shelf, $\varphi_{a}:X\to X$ a bijective map, for every $a\in X$, set $a\cdot b:=\varphi^{-1}_{a}\left(b\right)$ and $a:b:=\varphi^{-1}_{a}\left(a\triangleright b\right)$, for all $a,b\in X$, one can check that the map $r$ in (2.4) is a solution if and only if $\left(X,\,\cdot,\,:\right)$ is a q-cycle set, namely the structure introduced by Rump in [38]. In this way, we obtain a connection between 2.15 and the result provided by Rump in [38, Proposition 1] on the correspondence between left non-degenerate solutions on a set $X$ and q-cycle sets on the same set $X$ (see also [11, p. 2]).
@@
Set $a\ast b:=\varphi^{-1}_{a}\left(b\right)$, for all $a,b\in X$, it is easy to check that the map $r$ in 2.20 satisfies (bI) if and only if $\left(a\ast b\right)\ast\left(a\ast c\right)=\left(b\ast b\right)\ast\left(b\ast c\right)$, for all $a,b,c\in X$, i.e, the structure $\left(X,\ast\right)$ is a twisted Ward left quasigroup. Hence, 2.20 coincides with the result provided by Stanovský and Vojtěchovský in [40, Proposition 2.2] in which they establish a correspondence between left non-degenerate idempotent solutions on a set $X$ and twisted Ward left quasigroups on the same set $X$.
@@
The description of left non-degenerate solutions in terms of shelves automorphisms in 2.15 allows for explicitly obtaining the Drinfel’d twists that realize the conjugation between a left-non degenerate solution and its derived. We note that a similar connection between derived solutions/shelves and left non-degenerate solutions also appears in [38, p. 145], even if not explicitly made.
@@
for all $a,b\in X$, is a bijective and non-degenerate solution on $X$. Let us note that the binary operation given by $a\circ b=a+f\left(k\right)+b$, for all $a,b\in X$, endows $X$ of a structure of group with identity $-f\left(k\right)$ and, for every $a\in X$, the inverse of $a$ is $a^{-}=-f\left(k\right)-a-f\left(k\right)$. Moreover, the structure $\left(X,+,\circ\right)$ is a (singular) near brace (that is similar to [19, Example 2.10]).
@@ 3.1. Yang-Baxter algebras
We note that, from now on, whenever we say algebra we mean a binary algebra in a universal algebra sense, c.f. [5].
@@
If $(X,m)$ be a Yang-Baxter algebra for some solution $r$ and $\varphi:X\times X\to X\times X$ is a $D$-isomorphism, then $(X,m\varphi)$ is a Yang-Baxter algebra for a solution $\varphi^{-1}r\varphi.$
@@
Let $(X,r)$ be a left non-degenerate solution and $(X,\triangleright_{r})$ the shelf associated to $r$. Then, if $x\in X$, the binary operation normal-$\bullet$ on $X$ defined by
@@
[…]
  1. (1)
    $a\cdot b=b\cdot a,$ $\leavevmode\nobreak\ a\mapsto f:=\sum_{a\in X}f_{a}e_{a,a},$ $f_{a}\in\mathbb{C}.$
  2. (2)
    $a\cdot b=\sigma_{a}(b)\cdot\tau_{b}(a),$ $\leavevmode\nobreak\ a\mapsto w_{a}:=\sum_{c\in X}e_{\sigma_{a}(c),c}.$ Indeed, using the fundamental condition from the braid equation, $\sigma_{a}(\sigma_{b}(c))=\sigma_{\sigma_{a}(b)}(\sigma_{\tau_{b}(a)}(c))$ and the Yang-Baxter algebra, it follows that, $w_{a}w_{b}=w_{\sigma_{a}(b)}w_{\tau_{b}(a)}$
  3. (3)
    $a\cdot b=b\cdot(b\triangleright a),$ $\leavevmode\nobreak\ a\mapsto q_{a}:=\sum_{c\in X}e_{c,a\triangleright c}.$ Indeed using the definition $\leavevmode\nobreak\ a\cdot b:=b\triangleright(a\triangleright x),$ and the Yang-Baxter algebra we obtain, $q_{a}q_{b}=q_{b}q_{b\triangleright a}.$
@@
The superscript $p$ stands for parametric; in our case we consider functions of three fixed parameters $z_{1,2},\ \xi\in X,$ then $f^{p}:=f(z_{1},z_{2},z)$ for any function $f$ of the parameters.
Let $(X,+,\circ)$ be a skew brace and $r^{z},\ r^{*p}:X\times X\to X\times X$ are solutions of the braid equation, such that $r^{p}(x,y)=(\sigma^{z}_{p}(y),\tau^{z}_{y}(x)),$ $r^{*p}(x,y)=(\hat{\sigma}^{p}_{x}(y),\hat{\tau}^{p}_{y}(x))$ [19],
@@
$(2)\quad\hat{\sigma}_{a}^{z}(b)=a\circ b\circ z_{1}-a\circ z+z_{2}$ (3.7)
@@
3.2. Generalised affine shelves and Yang-Baxter algebras

### v3

@@ Abstract.
We present connections between left non-degenerate solutions of the set-theoretic braid equation and left shelves using Drinfel’d homomorphisms. We generalize the notion of affine quandle, by using heap endomorphisms and metahomomorphisms, and identify the underlying Yang-Baxter algebra for solutions of the braid equation associated to a given quandle. We introduce the notion of the pre-Lie skew brace and identify certain affine quandles that give rise to pre-Lie skew braces. Generalisations of the braiding of a group, associated to set-theoretic solutions of the braid equation are also presented. These generalized structures encode part of the underlying Hopf algebra. We then introduce the quasi-triangular (quasi) Hopf algebras and universal ${\cal R}$-matrices for rack and set-theoretic algebras. Generic set-theoretic solutions coming from heap endomorphisms are also identified.
@@ 1. Introduction
The Yang-Baxter equation is a key mathematical object in the theory of quantum integrable models and solvable statistical systems [32], as well as in the formulation of quantum groups [32, 43, 44, 27]. It was first introduced in [65] as the main mathematical tool for the investigation of quantum systems with many particle interactions, and in [6] for the study of a two-dimensional solvable statistical model known as the anisotropic Heisenberg magnet or the XYZ model. In early 90s Drinfeld [26] suggested the idea of set-theoretic solutions to the Yang-Baxter equation and since then, set-theoretic solutions have been extensively investigated primarily by means of representations of the braid group. It is worth noting that set-theoretic solutions and Yang-Baxter maps have been also studied within the context of classical discrete integrable systems connected also to the notion of Darboux-Bäcklund transformation in the Lax pair formulation [1, 52, 64]. In classical integrable systems usually a Poisson structure exists associated to a classical $r$-matrix, which is a solution of the classical Yang-Baxter equation [31]. Various connections between the set-theoretic Yang-Baxter equation and geometric crystals [8, 30], or soliton cellular automatons [39, 62] have been also demonstrated. The theory of braces was established around 2005, when Wolfgang Rump developed an algebraic structure that generalizes radical rings called a brace [54, 55] to describe all finite involutive set-theoretic solutions of the braid (and hence the Yang-Baxter) equation, whereas skew-braces were developed by Guarnieri and Vendramin to describe non-involutive solutions [38].
The investigation of set-theoretic solutions of the braid (and Yang-Baxter) equation and the associated algebraic structures is an emerging research area that has been particularly fruitful, and a significant number of related studies has been produced over the past few years (see for instance [7][19] [33][35], [41, 57, 58]). In [20][24] key links between set-theoretic solutions and quantum integrable systems and the associated quantum algebras were uncovered. Moreover, interesting correspondences between Yang-Baxter algebras and pre-Lie algebras found in [25], whereas connections between braces and pre-Lie rings established in [59]. In general, the theory of the set-theoretic Yang-Baxter equation turns out to have intriguing interlinks with numerous research areas, for example with group theory (Garside groups, regular subgroups, factorized groups, see for example [7, 41, 42]), algebraic number theory, Hopf-Galois extensions [7, 57], non-commutative rings [55], Knot theory [45][48], Hopf algebras, quantum groups [29], universal algebras, groupoids [53], semi-braces [15], trusses and heaps [11], pointed Hopf algebras, Yetter-Drinfield modules and Nichols algebras [3].
We now introduce the notion of the set-theoretic braid equation. Following [26], given a set $X$, a map $r:X\times X\to X\times X$ is said to be a set-theoretic solution of the Yang-Baxter equation, if $r$ satisfies the braid identity
@@
It is also worth recalling the connection between the set-theoretic braid equation (1.1) and the set-theoretic Yang-Baxter equation (see also [43, 44]). We introduce the map $R:X\times X\rightarrow X\times X,$ such that $R=r\pi,$ where $\pi:X\times X\rightarrow X\times X$ is the flip map: $\pi(x,y)=(y,x).$ Hence, $R(y,x)=(\sigma_{x}(y),\tau_{y}(x)),$ and it satisfies the set-theoretic Yang-Baxter equation:
$R_{12}\ R_{13}\ R_{23}=R_{23}\ R_{13}\ R_{12},$ (1.2)
where we denote $R_{12}(y,x,w)=(\sigma_{x}(y),\tau_{y}(x),w),$ $R_{23}(w,y,x)=(w,\sigma_{x}(y),\tau_{y}(x))$ and
$R_{13}(y,w,x)=(\sigma_{x}(y),w,\tau_{y}(x)).$
@@
The key outcomes of this investigation are essentially summarized in Theorem 2.15, where it is shown that every left non-degenerate solution can be obtained from a shelf solution by finding special shelf automorphisms, and in Sections 3-5. In Sections 3-5 we generalize the notion of affine quandle, by using heap endomorphisms and metahomomorphisms, and identify the Yang-Baxter algebra for solutions of the braid equation associated to a given rack/quandle. Motivated by the notion of pre-Lie algebras (also studied under the name chronological algebras) [2, 36, 63] (see also [5, 51] for a recent reviews) we introduce a novel algebraic structure called pre-Lie skew brace to describe the underlying algebraic structures associated to certain set-theoretic solutions of the braid equation. In fact, we identify families of affine quandles that produce pre-Lie skew braces. We then consider the linearized version of the braid equation and we introduce the notion of the quasi-triangular Hopf algebras associated to set-theoretic solutions; the rack type universal ${\cal R}-$matrix is also derived. By identifying a suitable admissible Drinfel’d twist [27, 28] we are able to extract the general set-theoretic universal ${\cal R}-$matrix. Our findings on universal admissible Drinfel’d twists are consistent with Theorem 2.15.
@@
More precisely, the outline of the paper is as follows. We start with Section 2 which is divided into two subsections. The first subsection contains preliminaries on shelves. In the second subsection, in 2.9, we introduce the notion of a Drinfel’d homomorphism between set-theoretic solutions of the Yang-Baxter equation, based on the well known idea of Drinfel’d twists. When one considers a Drinfel’d automorphism, one gets a set-theoretic analogue of admissible twists of quasitriangular Hopf-algebras. We show that every left non-degenerate solution is Drinfel’d isomorphic to a solution given by a shelf, see Lemma 2.12. In fact, by 2.15, every left non-degenerate solution can be acquired from a shelf solution by finding special automorphisms of the shelf itself, which we name twists, see 2.14. In particular, bijective solutions correspond to racks. We show that two homomorphic shelves with twists have D-homomorphic solutions, and bijectivity of a homomorphism of shelves implies D-isomorphism of solutions. We conclude Section 2 with examples of solutions given by shelves and twists on them.
@@
Section 3 is divided into three subsections. In the first subsection, we present a definition of a Yang-Baxter algebra $(X,m)$ associated with a given solution $(X,r)$ as a binary operation $m:X\times X\to X$ such that $mr=m,$ i.e., by the Yang-Baxter algebra in 3.1; specifically, we understand an algebra in the sense of a universal algebra that is invariant under the entanglement of a particular set-theoretic solution of the Yang-Baxter equation. We show that for any set endowed with a left non-degenerate solution, the Yang-Baxter algebra exists, see Lemma 3.4. We conclude the first subsection with examples of Yang-Baxter algebras for solutions given by skew braces and near braces. In the second subsection, we introduce three generalisations of affine quandles to the non-abelian case and study their Yang-Baxter algebras, see examples 3.12,3.13 and 3.14. In particular, endofunctions called metahomomorphisms and heap endomorphisms allow us to define such generalized quandles. Moreover, if the underlying group of those quandles is abelian and metahomomorphisms are group endomorphisms, all the examples collapse to the definition of an affine quandle. In addition, every heap endomorphism gives rise to quandle given by a metahomomorphism, see Lemma 3.16; nevertheless, the Yang-Baxter algebra operation obtained by a heap endomorphism is not necessarily given by a metahomomorphism, see 3.11 and Lemma 3.20. In the third and last subsection, we introduce a notion of the right pre-Lie skew brace, see 3.23. In particular, every pre-Lie ring is a right Pre-Lie brace. Next, we show that Yang-Baxter algebras from examples 3.13 and 3.14 admit a pre-Lie skew brace structure, see 3.28. If an underlying group of a pre-Lie skew brace is abelian, one can acquire a pre-Lie ring. We conclude the subsection with 3.34 in which we show that Lie rings associated with a Yang-Baxter algebra given by affine quandle have a zero Lie bracket.
@@
In Section 4, we present a generalized version of the braiding of a group, with some examples that have already appeared in [24, 23], see 4.1. We generalize it beyond the group structure to be compatible with Yang-Baxter algebras. We conclude the section with examples of such braidings, see examples 4.6,4.7,4.8 and 4.9. This section serves as a precursor of the last section given that the generalized braided structures encapsulate part of the associated underlying Hopf algebras, which is introduced next.
@@
In Section 5, which is divided in two subsections, we shift our focus from set-theoretic solutions to linearized versions and Hopf algebras. Specifically, in the first subsection, introduce the Yang-Baxter algebras of rack/quandle type and set-theoretic solutions as quasitriangular (quasi)-Hopf algebras [27, 28]. Relevant earlier works in [21, 22, 23] also in [29] and [3] in connection with pointed Hopf algebras, and the bialgebras associated to racks are studied in [19, 47, 4]. We first introduce the quandle Hopf algebra and we then systematically construct the associated universal ${\cal R}$-matrix. In the second subsection, we suitably extend the quandle algebra and present the set-theoretic Yang-Baxter algebra, which is also a Hopf algebra. By means of a a suitable admissible universal Drinfel’d twist we construct the universal-set theoretic ${\cal R}-$matrix. We note that the linearized version of the Yang-Baxter equation allows the investigation of the full Hopf algebras and not only the study of the subset consisting of group-like elements that naturally appear in the set-theoretic frame. Fundamental representations of the aforementioned algebras are also considered leading to the rack and general set-theoretic solutions of the Yang-Baxter equation. Generalized novel set-theoretic solutions coming from heap endomorphisms and being compatible with the universal set-theoretic Yang-Baxter algebra are also introduced.
@@ 2.1. Left shelves
Let $X$ be a non-empty set and  $\triangleright$  a binary operation on $X$. Then, the pair $\left(X,\,\triangleright\right)$ is said to be a left shelf if  $\triangleright$  is left self-distributive, namely, the identity
@@
If $\left(X,\,\triangleright\right)$ and $\left(Y,\,\blacktriangleright\right)$ are left shelves, a map $f:X\to Y$ is said to be a shelf homomorphism if $f\left(a\triangleright b\right)=f\left(a\right)\blacktriangleright f\left(b\right)$, for all $a,b\in X$, or, equivalently, $fL_{a}\left(b\right)=\bar{L}_{f\left(a\right)}f\left(b\right)$, for all $a,b\in X$, where $\bar{L}_{u}\left(v\right)=u\blacktriangleright v$, for all $u,v\in Y$. We denote the set of all endomorphisms of $\left(X,\,\triangleright\right)$  by $\mathrm{End}(X,\,\triangleright)$ and the set of all automorphisms of $\left(X,\,\triangleright\right)$  by $\mathrm{Aut}(X,\,\triangleright)$.
@@ 2.2. Left-non degenerate solutions and left shelves
The converse of the previous fact does not hold.
@@
Given a set $X$, a map $r:X\times X\to X\times X,$ such that $r(a,b)=(\sigma_{a}(b),\tau_{b}(a)),$ for all $a,b\in X,$ satisfies (1.1) if and only if the following three conditions are satisfied for all $a,b,c\in X$
@@
[…]
@@
As a consequence of 2.16 and [14, Corollary 6] we can state the following result.
@@
If $\left(X,\,\triangleright\right)$ is a left shelf, $\varphi_{a}:X\to X$ a bijective map, for every $a\in X$, set $a\cdot b:=\varphi^{-1}_{a}\left(b\right)$ and $a:b:=\varphi^{-1}_{a}\left(a\triangleright b\right)$, for all $a,b\in X$, one can check that the map $r$ in (2.4) is a solution if and only if $\left(X,\,\cdot,\,:\right)$ is a q-cycle set, namely the structure introduced by Rump in [56]. In this way, we obtain a connection between 2.15 and the result provided by Rump in [56, Proposition 1] on the correspondence between left non-degenerate solutions on a set $X$ and q-cycle sets on the same set $X$ (see also [14, p. 2]).
@@
Set $a\ast b:=\varphi^{-1}_{a}\left(b\right)$, for all $a,b\in X$, it is easy to check that the map $r$ in 2.20 satisfies (bI) if and only if $\left(a\ast b\right)\ast\left(a\ast c\right)=\left(b\ast b\right)\ast\left(b\ast c\right)$, for all $a,b,c\in X$, i.e, the structure $\left(X,\ast\right)$ is a twisted Ward left quasigroup. Hence, 2.20 coincides with the result provided by Stanovský and Vojtěchovský in [61, Proposition 2.2] in which they establish a correspondence between left non-degenerate idempotent solutions on a set $X$ and twisted Ward left quasigroups on the same set $X$.
@@
The description of left non-degenerate solutions in terms of automorphisms of shelves in 2.15 allows for explicitly obtaining the Drinfel’d twists that realize the conjugation between a left-non degenerate solution and its derived. We note that a similar connection between derived solutions/shelves and left non-degenerate solutions also appears in [56, p. 145], even if not explicitly made.
@@
for all $a,b\in X$, is a bijective and non-degenerate solution on $X$. Let us note that the binary operation given by $a\circ b=a+f\left(k\right)+b$, for all $a,b\in X$, endows $X$ of a structure of group with identity $-f\left(k\right)$ and, for every $a\in X$, the inverse of $a$ is $a^{-}=-f\left(k\right)-a-f\left(k\right)$. Moreover, the structure $\left(X,+,\circ\right)$ is a (singular) near brace (that is similar to [24, Example 2.10]).
@@ 3.1. Yang-Baxter algebras
We note that in this section, whenever we say algebra we mean a binary algebra in a universal algebra sense, c.f. [9].
@@
If $(X,m)$ is a Yang-Baxter algebra for some solution $r$ and $\varphi:X\times X\to X\times X$ is a $D$-isomorphism, then $(X,m\varphi)$ is a Yang-Baxter algebra for a solution $\varphi^{-1}r\varphi.$
@@
Let $(X,r)$ be a left non-degenerate solution and $(X,\triangleright_{r})$ the shelf associated to $r$. Then, if $x\in X$, the binary operation $\bullet$ on $X$ defined by
@@
The superscript $p$ stands for parametric; in our case we consider functions of three fixed parameters $z_{1,2},\ z\in X,$ then $f^{p}:=f(z_{1},z_{2},z)$ for any function $f$ of the parameters.
Let $(X,+,\circ)$ be a skew brace and $r^{p},\ r^{*p}:X\times X\to X\times X$ are solutions of the braid equation, such that $r^{p}(x,y)=(\sigma^{p}_{x}(y),\tau^{p}_{y}(x)),$ $r^{*p}(x,y)=(\hat{\sigma}^{p}_{x}(y),\hat{\tau}^{p}_{y}(x))$ [24],
@@
$(2)\quad\hat{\sigma}_{a}^{p}(b)=a\circ b\circ z_{1}-a\circ z+z_{2}$ (3.7)
@@
3.2. Generalized affine shelves and Yang-Baxter algebras